# Eccentricity distribution of extreme mass ratio inspirals

### v1

@@ I Introduction
Massive black holes (MBHs) with masses in the range of $10^{6}\,\text{-}\,10^{10}\,\mathrm{M_{\odot}}$ are commonly found at the cores of most galaxies [1, 2, 3]. Lighter MBH candidates with masses of $10^{4}\,\text{-}\,10^{5}\,\mathrm{M_{\odot}}$ have also been identified in dwarf galaxies [4]. These MBHs typically reside in dense stellar environments known as nuclear star clusters [5, 6, 7, 8, 9, 10], where frequent dynamical interactions occur [11, 12, 13]. Evidence of such interactions arises in the electromagnetic spectrum due to extreme phenomena like tidal disruption events involving stars and MBHs (see Refs. [14, 15] for recent reviews).
@@
The upcoming Laser Interferometer Space Antenna (LISA, see Refs. [16, 17]) will open a new observational window onto galactic nuclear interactions by detecting gravitational waves (GWs, see Ref. [18]) from compact objects, likely stellar-mass black holes (BHs), spiraling into MBHs. These systems, known as extreme mass ratio inspirals (EMRIs, see Refs. [19, 20, 21] for reviews) due to typical mass ratios in the range $10^{-5}\,\text{-}\,10^{-3}$, offer a unique probe of strong-field gravity.
@@
EMRI detections will yield insights in astrophysics [22], cosmology [23], fundamental physics [24], and waveform modeling [25]. Indeed, EMRIs are predicted to spend years in LISA band, accumulating high signal-to-noise ratios over around $10^{4}\,\text{-}\,10^{5}$ orbital cycles [26]. This will enable extremely precise measurements of MBH masses and spins [27] for the otherwise quiescent MBH population not emitting in the electromagnetic spectrum [28, 29]. The long-lived GW signal also makes EMRIs sensitive to environmental effects [30]: the presence of matter near the MBH might cause slight phase shifts in the waveform, allowing LISA to test for accretion disks [31, 32, 30, 33] and dark matter structures [34].
@@
EMRIs have also been proposed as sources of quasi-periodic eruptions (QPEs, e.g. Ref. [35]), currently unexplained soft X-ray flares repeating over few hours to few days, coming from the nuclei of some galaxies [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. This could enable multi-messenger observations of EMRIs in the future, although current QPE detections seem to point to precursors of LISA detectable EMRIs orbiting at lower frequencies [47].
@@
To fully exploit EMRI detections with LISA, we need fast and accurate waveform models for them. This has motivated the development of the FastEMRIWaveforms (FEW, see Refs. [48, 49, 50, 51, 52]) package, which is widely adopted in the community. FEW can generate fully relativistic, adiabatic waveforms for equatorial orbits around Kerr MBHs in under a second, although this is currently limited to eccentricities below 0.9.
@@
Whether this is a limitation or not for EMRI scientific exploitation, it depends on the considered formation channel. For example, this is not expected to be an issue for EMRIs forming in accretion disks [53, 54, 55], where gas interactions circularize the orbit, or for EMRIs forming from the tidal separation of stellar binaries [56, 57] which also tend to be only moderately eccentric. However, other formation channels might yield highly eccentric EMRIs within the LISA band. For instance, the Lidov–Kozai mechanism [58, 59, 60], triggered by a secondary MBH, can drive large eccentricity oscillations on the orbit of the smaller BH [61, 62, 63, 64]. EMRIs may also form following supernova kicks [65, 66], again leading to high eccentricities.
@@
Although all of the aforementioned channels might be at work, significantly contributing to the LISA EMRI rates, this latter is expected to be dominated by what is considered the “standard” EMRI formation mechanism: two body relaxation in dense nuclei (see Ref. [20] and references therein). In short, objects orbiting an MBH in nuclear star clusters undergo frequent orbital deflections through weak two-body encounters with nearby bodies, which are themselves subject to the same interactions. This relaxation process is mainly driven by angular momentum exchanges, which randomly modify orbital eccentricities until one orbit reaches a sufficiently small pericentre for GWs emission to dominate over two-body relaxation, thus triggering the formation of an EMRI. In this work, we focus on this mechanism and we derive the expected EMRI eccentricity distribution for Schwarzschild MBHs to verify whether current waveform models meet the needs of what is considered to be the dominant EMRI formation channel.
@@
Specifically, we build upon the results of Mancieri et al. [67], who modeled EMRI formation using a Monte Carlo scheme to consistently account for perturbations acting on the 2.5 post-Newtonian (PN, see Ref. [68]) orbital motion of a single stellar-mass BH due to weak encounters with other objects surrounding the central MBH. That study focused on EMRI formation and thus stopped simulations before the final inspiral. Here, we take their endpoint binary configurations and evolve them to plunge using FEW, to then apply appropriate weights to recover the underlying astrophysical eccentricity distribution.
@@ II Methods
In this section, we describe the methods we use to extract the eccentricity distribution of the EMRI population generated by Mancieri et al. [67] at various stages of their inspiral. We first detail the evolution scheme we employ to integrate this EMRI population up to the final plunge, then the weighting procedure we follow to produce astrophysically consistent eccentricity distributions.
@@ II.1 EMRI trajectory integration
In the simulations by Mancieri et al. [67], EMRIs were identified, and their integration stopped, as soon as their GW emission timescale $t_{\mathrm{GW}}=10^{-3}\,t_{\mathrm{rlx}}$, where $t_{\mathrm{rlx}}$ is the angular momentum relaxation time of the orbit (see Fig. 1) [^1] In the case of $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$, a significant fraction of EMRIs were stopped at $t_{\mathrm{GW}}<t_{\mathrm{rlx}}=10\,\mathrm{Myr}$, before the condition $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ was reached. A similar correction was also applied to lower MBH masses, but it impacted a much smaller number of EMRIs.. From this point onward, it is safe to assume that the dynamics is completely decoupled from two-body relaxation and determined by GWs emission only.
@@
Figure 1: Two example EMRIs forming and plunging into a $10^{5}\,\mathrm{M_{\odot}}$ MBH. The early part of the integration (blue shaded tracks) is from Mancieri et al. [67], while the final phase (blue solid tracks) is presented in this work. The gray dotted line marks the separatrix, and the gray shaded region below it is the “relativistic” loss cone. The loss cone is usually delimited by the pericenter condition $r_{\mathrm{p}}=8R_{\mathrm{g}}$, while here the separatrix corresponds to $r_{\mathrm{p}}=4R_{\mathrm{g}}$ for $e\to 1$ (see Sect. II.2). The green dashed curve separates the region dominated by two-body relaxation (above) from that dominated by GWs emission (below). Black crosses mark the initial conditions from Mancieri et al. [67], as well as the initial semi-major axes $a_{\mathrm{i}}$ used in this work to weight EMRIs (see Sect. II.2). Black dots indicate the stopping points of runs in Mancieri et al. [67], which also serve as initial conditions in this work. Both an initially wide ($a_{\mathrm{i}}>a_{\mathrm{c}}$) and a classical ($a_{\mathrm{i}}<a_{\mathrm{c}}$) EMRI are shown (see Sect. III).
From the output of the simulations by Mancieri et al. [67], we extract the final 2PN orbital energy (not including the stellar potential) and angular momentum $L$ per unit of the reduced mass of the binary (the mass of the secondary is fixed to $10\,\mathrm{M_{\odot}}$ throughout this work). We then find the energy $E$ of the corresponding geodesic by adding $c^{2}$ to the 2PN estimate to account for the rest energy of the secondary [^2] A more accurate conversion would be $E_{\mathrm{GR}}=c^{2}\sqrt{2E_{\mathrm{Newt}}/c^{2}+1}$, but it expands to $E_{\mathrm{GR}}\approx E_{\mathrm{Newt}}+c^{2}$ for small $E_{\mathrm{Newt}}/c^{2}$. This is always the case for the highly eccentric orbits we are considering. [69]. $E$ and $L$ are then converted into the the initial $(p_{0},e_{0})$ that we wish to integrate up to plunge, where $p$ is the semi-latus rectum and $e$ is the eccentricity of the orbit. We verify that this conversion returns sensible values, compatible to approximate estimates obtained considering the final pericenter and apocenter passages of the original simulations in Mancieri et al. [67]. To perform the change of parameters we use the ELQ_to_pex function provided by the FEW package, which implements the $(E,L_{\mathrm{z}},Q)\to(p,e,x_{\mathrm{I}})$ map developed by Hughes [70]. Here $Q$ is the Carter constant [71], while $x_{\mathrm{I}}$ is the cosine of the maximum inclination angle between the orbital and the MBH spin planes (see e.g. Ref. [72]), measured in Boyer-Lindquist coordinates [73]. As we consider Schwarzschild MBHs in this work, $Q=0$ and $x_{\mathrm{I}}=1$. For the same reason, the angular momentum along the MBH spin direction $L_{\mathrm{z}}$ is simply the same as $L$.
@@
In turn, $r_{\mathrm{p}}$ and $r_{\mathrm{a}}$ are the two largest roots of the radial effective potential of the MBH, given $E$ and $L$ (e.g. Refs. [74, 75, 76, 77]).
To integrate ($p,e$) down to plunge from their initial values ($p_{0},e_{0}$), we employ the trajectory package within FEW. The latest version of FEW includes two main trajectory modules [51]:
@@
    KerrEccEqFlux. The module is based on numerical computation of time-averaged fluxes in energy and angular momentum lost by the system (both to infinity and through the event horizon of the primary) due to GWs radiation. The fluxes are computed from the Teukolsky equation [78] using BH perturbation theory techniques (e.g. Refs. [79, 80, 81]). The computation of the fluxes is highly expensive, and thus it is performed offline for some discrete values of the orbital parameters. During the integration, this trajectory module simply interpolates over the precomputed grid of fluxes. As a result, the module can only be applied to pairs $(p,e)$ within the precomputed range. In the case of Schwarzschild MBHs, for $p\gtrsim 16.8\,R_{\mathrm{g}}$, this limits the eccentricity to $e<0.9$, where $R_{\mathrm{g}}=GM_{\bullet}/c^{2}$ is the gravitational radius of the MBH. For lower values of the semi-latus rectum, the eccentricity is further limited, dropping to $e\lesssim 0.82$ for the smallest allowed $p$ (see Fig. 2).
@@
    PN5. The module is based on analytical expressions for the inspiraling force, valid to $\mathcal{O}\big{(}(v/c)^{11},e^{11}\big{)}$ order [79]. Here $v/c=\sqrt{R_{\mathrm{g}}/p}$ is the small parameter over which the PN expansion is performed. This module can be applied at any $(p,e)$, even though its accuracy decreases for small $p$ and large $e$.
@@
Our initial values $e_{0}$ (roughly ranging from 0.99 to 0.9997) always fall outside the domain of validity of the KerrEccEqFlux trajectory model. We therefore proceed with the integration in two steps: (i) we integrate with the PN5 module to bring $(p,e)$ inside the range of the precomputed flux grids; (ii) we switch to the KerrEccEqFlux module to get the system to the plunge. In some cases, the plunge occurs before it is possible to switch module, so that the whole integration is performed with the PN5 module only. This is always the case for EMRIs that plunge at $e_{\mathrm{pl}}\gtrsim 0.82$ (see Fig. 2). The plunging condition is identified by $p=p_{\mathrm{sep}}$, where $p_{\mathrm{sep}}$ is the semi-latus rectum of the innermost bound stable orbit, also known as separatrix. For a Schwarzschild MBH [^3] To mitigate numerical issues that might arise using the full separatrix expression for generic orbits around Kerr MBHs, we perform the integration enabling the enforce_schwarz_sep flag in the trajectory class. this is given by [82]
$p_{\mathrm{sep}}=6+2e\,.$ (2)
We repeat the integration for all initial conditions twice: (i) by directly integrating the orbital elements $(p,e)$, and (ii) by exploiting the constants of motion $(E,L)$. The latter choice is more numerically robust near the separatrix [70], although issues can arise for almost circular orbits, where small errors in the constants of motion can be amplified when transformed back to $(p,e)$ within FEW. For this work, we opted for the integration of $(E,L)$, as we found that both methods give almost indistinguishable results at the precision level we require.
Figure 2 shows a selection of the EMRI evolution track that we produce.

### v2

@@ I Introduction
Massive black holes (MBHs) with masses in the range of $10^{6}\,\text{-}\,10^{10}\,\mathrm{M_{\odot}}$ are commonly found at the cores of most galaxies [1995ARA&A..33..581K, 1998AJ....115.2285M, 2013ARA&A..51..511K]. Lighter MBH candidates with masses of $10^{4}\,\text{-}\,10^{5}\,\mathrm{M_{\odot}}$ have also been identified in dwarf galaxies [2020ARA&A..58..257G]. These MBHs typically reside in dense stellar environments known as nuclear star clusters [1997AJ....114.2366C, 2002AJ....123.1389B, 2006ApJS..165...57C, 2007A&A...469..125S, 2009A&A...502...91S, 2012ApJS..203....5T], where frequent dynamical interactions occur [2008gady.book.....B, 2013degn.book.....M, 2017ARA&A..55...17A]. Evidence of such interactions arises in the electromagnetic spectrum due to extreme phenomena like tidal disruption events involving stars and MBHs (see Refs. [2021ARA&A..59...21G, 2021TDE] for recent reviews).
@@
The upcoming Laser Interferometer Space Antenna (LISA, see Refs. [2017arXiv170200786A, 2024arXiv240207571C]) will open a new observational window onto galactic nuclear interactions by detecting gravitational waves (GWs, see Ref. [2007gwte.book.....M]) from compact objects, likely stellar-mass black holes (BHs), spiraling into MBHs. These systems, known as extreme mass ratio inspirals (EMRIs, see Refs. [2007CQGra..24R.113A, 2018LRR....21....4A, 2022hgwa.bookE..17A] for reviews) due to typical mass ratios in the range $10^{-5}\,\text{-}\,10^{-3}$, offer a unique probe of strong-field gravity.
@@
EMRI detections will yield insights in astrophysics [2023LRR....26....2A], cosmology [2023LRR....26....5A], fundamental physics [2022LRR....25....4A], and waveform modeling [2023arXiv231101300L]. Indeed, EMRIs are predicted to spend years in LISA band, accumulating high signal-to-noise ratios over around $10^{4}\,\text{-}\,10^{5}$ orbital cycles [2004PhRvD..69h2005B]. This will enable extremely precise measurements of MBH masses and spins [2017PhRvD..95j3012B] for the otherwise quiescent MBH population not emitting in the electromagnetic spectrum [2010PhRvD..81j4014G, 2019ApJ...883L..18G]. The long-lived GW signal also makes EMRIs sensitive to environmental effects [2014PhRvD..89j4059B]: the presence of matter near the MBH might cause slight phase shifts in the waveform, allowing LISA to test for accretion disks [2011PhRvD..84b4032K, 2011PhRvL.107q1103Y, 2014PhRvD..89j4059B, 2025PhRvD.111h4006D] and dark matter structures [2024PhRvL.133l1404D].
@@
EMRIs have also been proposed as sources of quasi-periodic eruptions (QPEs, e.g. Ref. [2023A&A...675A.100F]), currently unexplained soft X-ray flares repeating over few hours to few days, coming from the nuclei of some galaxies [2019Natur.573..381M, 2020A&A...636L...2G, 2021Natur.592..704A, 2024A&A...684A..64A, 2025arXiv250617138A, 2021ApJ...921L..40C, 2025ApJ...983L..39C, 2023A&A...675A.152Q, 2024Natur.634..804N, 2025MNRAS.540...30B, 2025NatAs...9..895H]. This could enable multi-messenger observations of EMRIs in the future, although current QPE detections seem to point to precursors of LISA detectable EMRIs orbiting at lower frequencies [2025arXiv250510488S].
@@
To fully exploit EMRI detections with LISA, we need fast and accurate waveform models for them. This has motivated the development of the FastEMRIWaveforms (FEW, see Refs. [2021PhRvL.126e1102C, 2021PhRvD.104f4047K, 2023arXiv230712585S, 2025arXiv250609470C, chapman_bird_2025_15630565]) package, which is widely adopted in the community. FEW can generate fully relativistic, adiabatic waveforms for equatorial orbits around Kerr MBHs in under a second, although this is currently limited to eccentricities below 0.9.
@@
Whether this is a limitation or not for EMRI scientific exploitation, it depends on the considered formation channel. For example, this is not expected to be an issue for EMRIs forming in accretion disks [2007MNRAS.374..515L, 2021PhRvD.103j3018P, 2023MNRAS.521.4522D], where gas interactions circularize the orbit, or for EMRIs forming from the tidal separation of stellar binaries [1988Natur.331..687H, 2005ApJ...631L.117M] which also tend to be only moderately eccentric. However, other formation channels might yield highly eccentric EMRIs within the LISA band. For instance, the Lidov–Kozai mechanism [1910AN....183..345V, 1962P&SS....9..719L, 1962AJ.....67..591K], triggered by a secondary MBH, can drive large eccentricity oscillations on the orbit of the smaller BH [2011ApJ...729...13C, 2014MNRAS.438..573B, 2022ApJ...927L..18N, 2022MNRAS.516.1959M]. EMRIs may also form following supernova kicks [2017MNRAS.469.1510B, 2019MNRAS.485.2125B], again leading to high eccentricities.
@@
Although all of the aforementioned channels might be at work, significantly contributing to the LISA EMRI rates, this latter is expected to be dominated by what is considered the “standard” EMRI formation mechanism: two body relaxation in dense nuclei (see Ref. [2018LRR....21....4A] and references therein). In short, objects orbiting an MBH in a nuclear star cluster undergo frequent orbital deflections through weak two-body encounters with other bodies, which are themselves subject to the same interactions. This relaxation process is mainly driven by angular momentum exchanges, which randomly modify the orbital eccentricity until the orbit reaches a sufficiently small pericenter for GWs emission to dominate over two-body relaxation, thus triggering the formation of an EMRI. In this work, we focus on this mechanism and we derive the expected EMRI eccentricity distribution for Schwarzschild MBHs, to verify whether current waveform models meet the needs of what is considered to be the dominant EMRI formation channel.
@@
Specifically, we build upon the results of 2025A&A...694A.272M, who modeled EMRI formation using a Monte Carlo scheme to consistently account for perturbations acting on the 2.5 post-Newtonian (PN, see Ref. [2024LRR....27....2S]) orbital motion of a single stellar-mass BH due to weak encounters with other objects surrounding the central MBH. That study focused on EMRI formation and thus stopped simulations before the final inspiral. Here, we take their endpoint binary configurations and evolve them to plunge using FEW, to then apply appropriate weights to recover the underlying astrophysical eccentricity distribution.
@@ II Methods
In this section, we describe the methods we use to extract the eccentricity distribution of the EMRI population generated by 2025A&A...694A.272M at various stages of their inspiral. We first detail the evolution scheme we employ to integrate this EMRI population up to the final plunge, then the weighting procedure we follow to produce astrophysically consistent eccentricity distributions.
@@ II.1 EMRI trajectory integration
In the simulations by 2025A&A...694A.272M, EMRIs were identified, and their integration stopped, as soon as their GW emission timescale $t_{\mathrm{GW}}=10^{-3}\,t_{\mathrm{rlx}}$, where $t_{\mathrm{rlx}}$ is the angular momentum relaxation time of the orbit (see Fig. 1) [^1] In the case of $M_{\bullet}=4\times 10^{6}\,\mathrm{M_{\odot}}$, a significant fraction of EMRIs were stopped at $t_{\mathrm{GW}}<t_{\mathrm{rlx}}=10\,\mathrm{Myr}$, before the condition $t_{\mathrm{GW}}=10^{-3}t_{\mathrm{rlx}}$ was reached. A similar correction was also applied to lower MBH masses, but it impacted a much smaller number of EMRIs.. From this point onward, it is safe to assume that the dynamics is completely decoupled from two-body relaxation and determined by GWs emission only.
@@
Figure 1: Two example EMRIs forming and plunging into a $10^{5}\,\mathrm{M_{\odot}}$ MBH. The early part of the integration (blue shaded tracks) is from 2025A&A...694A.272M, while the final phase (blue solid tracks) is presented in this work. The gray dotted line marks the separatrix, and the gray shaded region below it is the “relativistic” loss cone. The loss cone is usually delimited by the pericenter condition $r_{\mathrm{p}}=8R_{\mathrm{g}}$, while here the separatrix corresponds to $r_{\mathrm{p}}=4R_{\mathrm{g}}$ for $e\to 1$ (see Sect. II.2). The green dashed curve separates the region dominated by two-body relaxation (above) from that dominated by GWs emission (below). Black crosses mark the initial conditions from 2025A&A...694A.272M, as well as the initial semi-major axes $a_{\mathrm{i}}$ used in this work to weight EMRIs (see Sect. II.2). Black dots indicate the stopping points of runs in 2025A&A...694A.272M, which also serve as initial conditions in this work. Both an initially wide ($a_{\mathrm{i}}>a_{\mathrm{c}}$) and a classical ($a_{\mathrm{i}}<a_{\mathrm{c}}$) EMRI are shown (see Sect. III).
From the output of the simulations by 2025A&A...694A.272M, we extract the final 2PN orbital energy (not including the stellar potential) and angular momentum $L$ per unit of the reduced mass of the binary (the mass of the secondary is fixed to $10\,\mathrm{M_{\odot}}$ throughout this work). We then find the energy $E$ of the corresponding geodesic by adding $c^{2}$ to the 2PN estimate to account for the rest energy of the secondary [^2] A more accurate conversion would be $E_{\mathrm{GR}}=c^{2}\sqrt{2E_{\mathrm{Newt}}/c^{2}+1}$, but it expands to $E_{\mathrm{GR}}\approx E_{\mathrm{Newt}}+c^{2}$ for small $E_{\mathrm{Newt}}/c^{2}$. This is always the case for the highly eccentric orbits we are considering. [2017grav.book.....M]. $E$ and $L$ are then converted into the the initial $(p_{0},e_{0})$ that we wish to integrate up to plunge, where $p$ is the semi-latus rectum and $e$ is the eccentricity of the orbit. We verify that this conversion returns sensible values, compatible to approximate estimates obtained considering the final pericenter and apocenter passages of the original simulations in 2025A&A...694A.272M. To perform the change of parameters we use the ELQ_to_pex function provided by the FEW package, which implements the $(E,L_{\mathrm{z}},Q)\to(p,e,x_{\mathrm{I}})$ map developed by 2024PhRvD.109f4077H. Here $Q$ is the Carter constant [1968PhRv..174.1559C], while $x_{\mathrm{I}}$ is the cosine of the maximum inclination angle between the orbital and the MBH spin planes (see e.g. Ref. [2024arXiv241104955L]), measured in Boyer-Lindquist coordinates [1967JMP.....8..265B]. As we consider Schwarzschild MBHs in this work, $Q=0$ and $x_{\mathrm{I}}=1$. For the same reason, the angular momentum along the MBH spin direction $L_{\mathrm{z}}$ is simply the same as $L$.
@@
In turn, $r_{\mathrm{p}}$ and $r_{\mathrm{a}}$ are the two largest roots of the radial effective potential of the MBH, given $E$ and $L$ (e.g. Refs. [2002CQGra..19.2743S, 2002PhRvD..66d4002G, 2009CQGra..26m5002F, 2020CQGra..37n5007V]).
To integrate ($p,e$) down to plunge from their initial values ($p_{0},e_{0}$), we employ the trajectory package within FEW. The latest version of FEW includes two main trajectory modules [2025arXiv250609470C]:
@@
    KerrEccEqFlux. The module is based on numerical computation of time-averaged fluxes in energy and angular momentum lost by the system (both to infinity and through the event horizon of the primary) due to GWs radiation. The fluxes are computed from the Teukolsky equation [1973ApJ...185..635T] using BH perturbation theory techniques (e.g. Refs. [2020PhRvD.102f4005F, 2021PhRvD.103j4014H, 2021PhRvD.103j4045S]). The computation of the fluxes is highly expensive, and thus it is performed offline for some discrete values of the orbital parameters. During the integration, this trajectory module simply interpolates over the precomputed grid of fluxes. As a result, the module can only be applied to pairs $(p,e)$ within the precomputed range. In the case of Schwarzschild MBHs, for $p\gtrsim 16.8\,R_{\mathrm{g}}$, this limits the eccentricity to $e<0.9$, where $R_{\mathrm{g}}=GM_{\bullet}/c^{2}$ is the gravitational radius of the MBH. For lower values of the semi-latus rectum, the eccentricity is further limited, dropping to $e\lesssim 0.82$ for the smallest allowed $p$ (see Fig. 2).
@@
    PN5. The module is based on analytical expressions for the inspiraling force, valid to $\mathcal{O}\big((v/c)^{11},e^{11}\big)$ order [2020PhRvD.102f4005F]. Here $v/c=\sqrt{R_{\mathrm{g}}/p}$ is the small parameter over which the PN expansion is performed. This module can be applied at any $(p,e)$, even though its accuracy decreases for small $p$ and large $e$.
@@
Our initial values $e_{0}$ (roughly ranging from 0.99 to 0.9997) always fall outside the domain of validity of the KerrEccEqFlux trajectory model. We therefore proceed with the integration in two steps: (i) we integrate with the PN5 module to bring $(p,e)$ inside the range of the precomputed flux grids; (ii) we switch to the KerrEccEqFlux module to get the system to the plunge. In some cases, the plunge occurs before it is possible to switch module, so that the whole integration is performed with the PN5 module only. This is always the case for EMRIs that plunge at $e_{\mathrm{pl}}\gtrsim 0.82$ (see Fig. 2). The plunging condition is identified by $p=p_{\mathrm{sep}}$, where $p_{\mathrm{sep}}$ is the semi-latus rectum of the innermost bound stable orbit, also known as separatrix. For a Schwarzschild MBH [^3] To mitigate numerical issues that might arise using the full separatrix expression for generic orbits around Kerr MBHs, we perform the integration enabling the enforce_schwarz_sep flag in the trajectory class. In typical runs the integration is terminated once $p/R_{\mathrm{g}}-6-2e=2\times 10^{-3}$. In case the integration is performed with the PN5 module only, it instead stops at $p/R_{\mathrm{g}}-6-2e=10^{-1}$. This affects only few systems and only the high-eccentricity tail ($e_{\mathrm{pl}}\gtrsim 0.82$) of Fig. 5. this is given by [2020PhRvD.101f4007S]
$p_{\mathrm{sep}}=(6+2e)R_{\mathrm{g}}\,.$ (2)
We repeat the integration for all initial conditions twice: (i) by directly integrating the orbital elements $(p,e)$, and (ii) by exploiting the constants of motion $(E,L)$. The latter choice is more numerically robust near the separatrix [2024PhRvD.109f4077H], although issues can arise for almost circular orbits, where small errors in the constants of motion can be amplified when transformed back to $(p,e)$ within FEW. For this work, we opted for the integration of $(E,L)$, as we found that both methods give almost indistinguishable results at the precision level we require.
Figure 2 shows a selection of the EMRI evolution tracks that we produce.